# AppApp Readings

### v1

@@ Abstract
We study doping and temperature dependencies of temporal and spatial spin stiffnesses of the Hubbard model within the mean field approach for incommensurate magnetic order. We show that the frequency dependence of temporal spin stiffness is crucial to obtain small values of correlation length, comparable to those observed in cuprates. Using the obtained spin stiffnesses, we obtain the temperature and doping dependence of correlation length within the large-$N$ limit of the respective nonlinear sigma model. In agreement with the experimental data on La2-xSrxCuO4 we obtain short range magnetic order with relatively small correlation length at $0.1\lesssim x\lesssim 0.2$, and magnetically ordered ground state in the narrow doping region $0.05\lesssim x\lesssim 0.1$. The latter state may correspond to the spin-frosen state, observed in the experimental data on La2-xSrxCuO4.
[…]
@@
To describe the low-energy spin excitations of magnetic systems the nonlinear sigma model (NLSM) can be used, see, e.g., Ref. Auerbach, for a review. The corresponding excitations in this model are parameterized by the temporal and spatial spin stiffnesses, which can be determined from the microscopic analysis. For insulating magnets, the classical version of the $O(3)$ non-linear sigma model was derived in the continuum limit of the classical Heisenberg model, assuming commensurate antiferromagnetic order ZinnJustin ; NelsonPelkovits . Later this derivation was also generalized to the quantum case Haldane ; CHN ; CSY . The nonlinear sigma models in the $O(3)/O(2)$ (Refs. DombreRead, ; Mouhanna, ) and $CP^{N-1}$ (Refs. CPN, ; CSS, ; Azaria, ) manifolds were considered in a continuum limit of frustrated quantum antiferromagnets with spin spiral ground state.
@@
The situation in itinerant systems is more involved. The derivation of the NLSM, which also allows obtaining the respective spin stiffnesses, was first proposed in Refs. Weng, ; Schulz, ; Dupuis, ; Dupuis1, ; Sachdev, ; Sachdev1, using Hubbard-Stratonovich transformation. This derivation introduced in the action the effective fluctuating field, corresponding to the order parameter, which is then fixed at its mean field value. More recently, using the SU(2)-symmetric gauge theory, which does not introduce symmetry breaking terms or condensation of gauge fields in the action, was proposed for derivation of the nonlinear sigma model and obtaining respective spin stiffnesses Bonetti ; BonettiThesis ; BonettiMetzner ; BonettiErratum ; OurSecond . In this approach the symmetry is broken only at the level of dressed single- and two-particle Green’s functions via spin-asymmetric self-energy and the resulting renormalized interaction vertices. The respective Ward identities for the latter approach were derived in Refs. Bonetti, ; BonettiThesis, ; BonettiErratum, ; OurSecond, . It was emphasized in Ref. OurSecond, that the temporal spin stiffness is in general dynamic, and possesses strong frequency dependence away from half filling. This frequency dependence becomes more pronounced with increase of doping.
@@
In the present paper, we consider the doping- and temperature dependence of spatial and temporal spin stiffnesses of the non-linear sigma model, obtained within the mean field approach for incommensurate magnetic order of the Hubbard model. While previous consideration BonettiMetznerStatic ; BonettiNew was restricted to the static limit, we uncover the role of the dynamic effects in temporal spin stiffness. We emphasize that strong frequency dependence, which occurs already in the low-energy regime, corresponds to consideration of gapped magnetic modes, which are absent in the standard spin-wave approaches, but were discussed recently within the bond-operator technique Syromyatnikov . Using the obtained results, we calculate the temperature- and doping dependence of correlation length and show that the gapped modes are crucial to obtain qualitative agreement with the experimental data on high-$T_{c}$ compounds.
Model and magnetic correlations. We consider the Hubbard model on the square lattice
$H=-\sum_{i,j,\sigma}t_{ij}{c_{i\sigma}^{\dagger}c_{j\sigma}}+U\sum_{i}{n_{i\uparrow}n_{i\downarrow}},$ (1)
with hopping $t_{ij}=t$ between nearest neighbors (which is used as a unit of energy) and $t_{ij}=-t^{\prime}$ for next-nearest neighbors.
@@
We consider spiral magnetic order in the $x,z$ plane. For the reasons described below, we consider mainly the out of plane susceptibility, which corresponds to the response $\chi^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}=\bar{\chi}^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}={-}\langle\langle S^{y}_{q}|S^{y}_{-q^{\prime}}\rangle\rangle$ (the bar denotes the local spin reference frame). This susceptibility depends on temporal and spatial spin stiffnesses $\chi_{\rm{op},\omega_{n}}$,$\rho_{\rm{op},xx/yy}$ as parameters, which can be determined from microscopic analysis, and characterizing the low energy magnetic excitations Bonetti ; BonettiErratum ; BonettiThesis ; BonettiMetzner ; OurSecond ,
@@
where $m$ is the long-range order parameter. Short-range magnetic order is characterized by spin gap $\Delta$, determining the correlation length $\xi$.
To obtain correlation length, we consider the leading order of $1/N$ expansion in the massive CPN-1 modelCPN ; CSS ; BonettiMetzner . In this case, only the out-of-plane mode of magnetic excitations contributes to the correlation length, describing low-lying magnetic excitations and short-range magnetic fluctuations, in the large number of components $N$ limit. The value of an energy gap $\Delta$ is determined in this approach by the respective sum rule CSS
@@
where the integration is carried over sufficiently small momenta $\mathbf{q}$ (the chosen integration region is specified in Supplemental MaterialSM ), the factor 1/2 in the right-hand side accounts for the presence of two non-identical magnetic wave vectors ±$\pm{\bf Q}$ in the Brillouin zone. In the polar coordinates this sum rule results in an equation for $\Delta$
@@
[…]
There are two distinct regimes of spin gap temperature dependence with lowering temperature. If we assume for $T\to 0$ that $\Delta\to 0$ than $\kappa(T)\to\kappa(0)\equiv\kappa>0$ and the equation for the spin gap simplifies to
@@
In this case, the role of dynamic quantum fluctuations is only in the renormalization of spatial spin stiffness $\bar{\rho}\to\kappa\bar{\rho}$ in the exponent, similarly to the previous analysis of $O(N)/O(N-1)$, Ref. CSY, and massive $CP^{N-1}$, Ref. CSS, nonlinear sigma models. In this regime according to the equation (7) we define the crossover temperature $T^{*}=2\pi\kappa\bar{\rho}$ to the quasi-long range order state (denoted also as the renormalized classical regime CHN ), such that at $T<T^{*}$ the correlation length gets exponentially large.
@@
Mean field calculations of spin stiffnesses. For calculations, we use the standard mean-field approach; the spin stiffnesses are expressed through the spin- and current correlation functions, see Refs. Bonetti, ; BonettiMetzner, ; BonettiThesis, ; BonettiErratum, ; OurSecond, and Supplemental Material SM . We choose $t$ as a unit of energy, $t^{\prime}/t=0.16$, which is close to that for La2CuO4, and the relatively weak interaction strength $U/t=2.5$, since for stronger interaction, the incommensurate state in the mean field approach is unstable at low dopings. In view of the used mean field approach, this interaction should be, however, considered as an effective screened interaction, which is typically much smaller than its bare value, cf. Refs. KataninYamase, ; IrkhinIgoshev, . The considered interaction also coincides with that used previously in Refs. BonettiMetznerStatic, ; BonettiThesis, ; it is also close to that obtained for not too small doping range in Ref. BonettiMetzner, .
@@
The obtained frequency dependence of the temporal spin stiffness (see Fig. 1a) is strong already for $\omega_{n}\ll 1$ (in units of $t$), where the NLSM is applicable. While the static limit varies only weakly with electron concentration $n$, the decrease of the high-frequency asymptote (which is also almost temperature-independent) corresponds to the strengthening of quantum fluctuations with lowering $n$. These quantum fluctuations are related to the finite spectral weight at finite energies. In Fig. 1b we show the resulting frequency dependence of the out-of-plane susceptibility at the real frequency axis, obtained by analytic continuation. At and close to half filling we find the peak at low frequencies corresponding to a vanishing or small spin gap. However, apart from that, we observe generally two other peaks. One can see that the position of the lower of the two peaks coincides with the low-energy edge of the inter band spectral function $A_{{\mathbf{Q}},\rm ib}(\omega)=\sum_{\mathbf{k}}\delta(E^{+}_{\mathbf{k}+\mathbf{Q}}-E^{-}_{\mathbf{k}}-\omega)(f(E^{-}_{\mathbf{k}})-f(E^{+}_{\mathbf{k}+\mathbf{Q}}))$, shown in Fig. 1, where $E_{\mathbf{k}}^{\pm}$ are the electronic energy bands. The low-energy edge of $A_{{\mathbf{Q}},\rm ib}$ is determined by an indirect gap in the electronic spectrum. This peak can be therefore identified with the Higgs mode Sachdev1 ; Sachdev2 ; Sachdev ; SachdevHiggs1 . On approaching the paramagnetic mean-field solution (i.e. at substantial doping levels) this peak shifts to very low energies, and therefore, becomes suppressed by the spin gap. Therefore, at sufficiently large dopings, we observe a single peak in the spin spectral functions. These peaks in the spin spectral function remind us also of the gapped magnetic excitations, discussed recently for the square-lattice Heisenberg model within the bond-operator technique Syromyatnikov .
@@
[…]
@@
In summary, we have analyzed the effect of the frequency dependence of temporal spin stiffness on magnetic order in the doped Hubbard model. We find that this frequency dependence is crucial to obtain short range magnetic order with a sufficiently small correlation length. In particular, the gapped energy modes, one of which corresponds to the Higgs mode, carry substantial spectral weight, which yields the decrease of correlation length, determined by the sum rule. As a result, we find long-range magnetic order only in the narrow domain of dopings $x\simeq 0.1$, in agreement with the experimental data. We expect that this region of long-range order may be further shrunk by fluctuations, and/or partly unstable with respect to the non-uniform patterns, see, e.g., recent study of Ref. MetznerFiniteLattice, , which finds stripe instability in approximately the same doping range.
@@
Note added. When finalizing the paper, we have learned about the recent study of Ref. BonettiNew, , which considered the temperature- and concentration dependencies of spin stiffnesses and correlation length. However, only static temporal spin stiffness was considered in that study.
Acknowledgements. This study is supported by the Russian Science Foundation (Grant No. 24-12-00186).
@@
.1 q-integration in the sumrule (3) and the choice of the cutoff parameter

### v3

@@ Abstract
We study doping and temperature dependencies of temporal and spatial spin stiffnesses of the Hubbard model within the mean field approach for incommensurate magnetic order. We show that the frequency dependence of temporal spin stiffness within the considered mean field approach is crucial to obtain small values of correlation length, comparable to those observed in cuprates. Using the obtained spin stiffnesses, we obtain the temperature and doping dependence of correlation length within the large-$N$ limit of the respective nonlinear sigma model. In agreement with the experimental data on La2-xSrxCuO4 we obtain short range magnetic order with relatively small correlation length at $0.1\lesssim x\lesssim 0.2$, and magnetically ordered ground state in the narrow doping region $0.05\lesssim x\lesssim 0.1$. The latter state may correspond to the spin-frosen state, observed in the experimental data on La2-xSrxCuO4.
Magnetic properties of high-$T_{c}$ compounds remain an active research field, despite long time of their study. For the most studied high-$T_{c}$ compound La2-xSrxCuO4 the commensurate long-range magnetic order is destroyed already at small doping (see, e.g., Refs. Plakida, ; StaticOrd, ), while the short-range magnetic order with the wave vector $\mathbf{Q}=(\pi,\pi-\delta)$ is formed with doping ExpQ1 ; ExpQ2 ; ExpQ3 ; StaticOrd ; Plakida . Interestingly, despite the presence of strong electronic correlations, for not too small dopings x=1n>0.03x=1-n>0.03 the correlation length is quite moderate ExpInel1 ; ExpInel2 , $\xi\sim(2$-$5)a$, except for the doping $x\simeq 0.12$ (where stripe appearance was suggested). At the same time, the signs of static spin freezing and long-range magnetic order were observed at $x\simeq 0.12$ (see Refs. LRO_in_cuprate_Katano2000, ; LRO_in_cuprate_Khaykovich2002, ; LRO_in_cuprate_Khaykovich2003, ; LRO_in_cuprate_Chang2008, ; LRO_in_cuprate_Khaykovich2005, ; LRO_in_cuprate_Lake2002_Nature, ; NMR1, ; NMR2, ; StrMet, ). These observations show a strong change of magnetic properties with doping even in the moderately doped regime $x\gtrsim 0.1$. A possible relation of short-range magnetic order to pseudogap formation was suggested and actively discussed PseudogapSF ; PseudogapSad ; PseudogapOnufr ; PseudogapToschi ; PseudogapMag1 ; PaseudogapMag2 ; PseudogapMag3 ; Sachdev1 ; Sachdev2 ; BonettiMetzner .
@@
To describe the low-energy spin excitations of magnetic systems the nonlinear sigma model (NLSM) can be used, see, e.g., Ref. Auerbach, for a review. The corresponding excitations in this model are parameterized by the temporal and spatial spin stiffnesses, which can be determined from the microscopic analysis. For insulating magnets, the classical version of the $O(3)$ nonlinear sigma model was derived in the continuum limit of the classical Heisenberg model, assuming commensurate antiferromagnetic order ZinnJustin ; NelsonPelkovits . Later this derivation was also generalized to the quantum case Haldane ; CHN ; CSY . The nonlinear sigma models in the $O(3)/O(2)$ (Refs. DombreRead, ; Mouhanna, ) and $CP^{N-1}$ (Refs. CPN, ; CSS, ; Azaria, ) manifolds were considered in a continuum limit of frustrated quantum antiferromagnets with spin spiral ground state.
@@
The situation in itinerant systems is more involved. The derivation of the NLSM, which also allows for obtaining the respective spin stiffnesses, was first proposed in Refs. Weng, ; Schulz, ; Dupuis, ; Dupuis1, ; Sachdev, ; Sachdev1, using Hubbard-Stratonovich transformation. This derivation introduced in the action the effective fluctuating field, corresponding to the order parameter, which is then fixed at its mean field value. More recently, using the SU(2)-symmetric gauge theory, which does not introduce symmetry breaking terms or condensation of gauge fields in the action, was proposed for derivation of the nonlinear sigma model and obtaining respective spin stiffnesses Bonetti ; BonettiThesis ; BonettiMetzner ; BonettiErratum ; OurSecond . In this approach the symmetry is broken only at the level of dressed single- and two-particle Green’s functions via spin-asymmetric self-energy and the resulting renormalized interaction vertices. The broken spin symmetry in the so-called chargon sector allows one to describe suppression of spectral weight at the points of the Fermi surface close to $(\pi,0)$ and $(0,\pi)$ (see, e.g., Ref. BonettiMetzner, ). The respective Ward identities for the SU(2) symmetric approach were derived in Refs. Bonetti, ; BonettiThesis, ; BonettiErratum, ; OurSecond, . It was emphasized in Ref. OurSecond, that the temporal spin stiffness is in general dynamic, and possesses strong frequency dependence away from half filling. This frequency dependence becomes more pronounced with increase of doping.
@@
In the present paper, we consider the doping- and temperature dependence of spatial and temporal spin stiffnesses of the nonlinear sigma model, obtained within the mean field approach for incommensurate magnetic order of the Hubbard model. While previous consideration BonettiMetznerStatic ; BonettiNew was restricted to the static limit, we uncover the role of the dynamic effects in temporal spin stiffness. We emphasize that strong frequency dependence, which occurs already in the low-energy regime, corresponds to consideration of gapped magnetic modes, which are absent in the standard spin-wave approaches, but were discussed recently within the bond-operator technique Syromyatnikov . Using the obtained results, we calculate the temperature- and doping dependence of correlation length and show that the gapped modes are crucial to obtain qualitative agreement with the experimental data on high-$T_{c}$ compounds.
@@
Magnetic correlations. We consider magnetic correlations in a generic electronic system with spiral magnetic order in the $x,z$ plane. For the reasons described below, we consider mainly the out of plane magnetic susceptibility in Matsubara formalism $\chi^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}=\bar{\chi}^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}={\int d\tau e^{i\omega_{n}\tau}\langle S_{\mathbf{q}}^{y}(\tau)S_{\mathbf{-}q}^{y}(0)\rangle}$ ($S^{y}_{\mathbf{q}}$ is the Fourier transform of $y$-component of the itinerant spin operator, the bar denotes the local spin frame, $\tau$ refers to imaginary time). This susceptibility depends on temporal and spatial spin stiffnesses $\chi_{\rm{op},\omega_{n}}$$\rho_{\rm{op},xx/yy}$ as parameters, which can be determined from microscopic analysis, and characterize the low energy magnetic excitations Bonetti ; BonettiErratum ; BonettiThesis ; BonettiMetzner ; OurSecond ,
@@
where $m$ is the mean-field long-range order parameter. The actual short-range magnetic order is characterized by spin gap $\Delta$, determining the correlation length $\xi$.
To obtain correlation length, we consider the leading order of $1/N$ expansion in the massive CPN-1 model CPN ; CSS ; BonettiMetzner ; SM . In this case, only the out-of-plane mode of magnetic excitations contributes to the correlation length, describing magnetic excitations in the large number of components $N$ limit. The value of an energy gap $\Delta$ is determined in this approach by the respective sum rule CSS ; SM
@@
where the integration is carried over sufficiently small momenta $\mathbf{q}$ (the chosen integration region is specified in Supplemental MaterialSM ), the factor 1/2 in the right-hand side accounts for the presence of two non-identical magnetic wave vectors $\pm{\bf Q}$ in the Brillouin zone. In the polar coordinates the equation for $\Delta$ reads
@@
where we characterize the spin gap by the averaged spatial spin stiffness $\bar{\rho}=(\rho_{\rm{op},x}\rho_{\rm{op},y})^{1/2}$, since spin stiffnesses $\rho_{x,y}$ are close to each other, and perform analytic continuation of the susceptibility $\chi_{{\rm op},\omega}$ to the whole imaginary axis, $\bar{q}_{m}^{2}=\exp(\int d\phi\ln q_{m}(\phi))/\pi$. Note that in general spatial spin stiffness $\bar{\rho}$ is temperature-dependent. This fact is of extreme importance at intermediate temperatures due to the proximity to phase transitions where spatial spin stiffness vanishes. Apart from this effect, temperature dependence of $\kappa(T)$ comes from two different sources: lower frequency integration bound $\pi T$ and temperature dependence of the energy gap $\Delta(T)$. Both significantly reduce the role of quantum fluctuations at finite temperatures. The corresponding correlation length can be estimated as $\xi(T)=\sqrt{\bar{\rho}/\Delta(T)}$.
There are two distinct regimes of spin gap temperature dependence with lowering temperature. If we assume that $\Delta\to 0$ for $T\to 0$ than $\kappa(T)\to\kappa(0)\equiv\kappa>0$ and the equation for the spin gap simplifies to
@@
In this case, the role of dynamic quantum fluctuations is only in the renormalization of spatial spin stiffness $\bar{\rho}\to\kappa\bar{\rho}$ in the exponent, similarly to the previous analysis of O(N)/O(N1)O(N)/O(N-1), Ref. CSY, and massive $CP^{N-1}$, Ref. CSS, nonlinear sigma models. In this regime according to the equation (6) we define the crossover temperature $T^{*}=2\pi\kappa\bar{\rho}$ to the quasi-long range order state (denoted also as the renormalized classical regime CHN ), such that at $T<T^{*}$ the correlation length gets exponentially large.
@@
Spin stiffnesses within the mean-field approach. For numerical analysis of magnetic excitations we consider Hubbard model on the square lattice as a prototypical model describing strong magnetic correlations,
$H=-\sum_{i,j,\sigma}t_{ij}{c_{i\sigma}^{\dagger}c_{j\sigma}}+U\sum_{i}{n_{i\uparrow}n_{i\downarrow}}.$ (8)
We suppose the hopping $t_{ij}=t$ between nearest neighbors (which is used as a unit of energy) and $t_{ij}=-t^{\prime}$ for next-nearest neighbors and employ a mean‑field approximation, which leads to a gapped electronic spectrum with dispersion $E^{\pm}_{\mathbf{k}}$. The magnetization $m$ is obtained via solution of the respective mean-field equations; minimization of the free energy allows us to find $\mathbf{Q}$} as a function of filling and temperature. Spin stiffnesses are expressed through the spin- and current correlation functions, subsequently calculated within the random phase approximation approach, see Refs. Bonetti, ; BonettiMetzner, ; BonettiThesis, ; BonettiErratum, ; OurSecond, and Supplemental Material SM for computational details.
@@
We choose next-nearest neighbor hopping $t^{\prime}/t=0.16$, which is close to that for La2CuO4, and the relatively weak interaction strength $U/t=2.5$, since for stronger interaction, the incommensurate state in the mean field approach is unstable at low dopings. In view of the used mean field approach, this interaction should be, however, considered as an effective screened interaction, which is typically much smaller than its bare value, cf. Refs. KataninYamase, ; IrkhinIgoshev, . The considered interaction also coincides with that used previously in Refs. BonettiMetznerStatic, ; BonettiThesis, ; it is also close to that obtained for not too small doping range in Ref. BonettiMetzner, .
@@
The obtained frequency dependence of the temporal spin stiffness (see Fig. 1a) is strong already for $\omega_{n}\ll 1$ (in units of $t$), where the NLSM is applicable. While the static limit varies only weakly with electron concentration $n$, the decrease of the high-frequency asymptote (which is also almost temperature-independent) corresponds to the strengthening of quantum fluctuations with lowering $n$. These quantum fluctuations are related to the finite spectral weight at finite energies. In Fig. 1b we show the resulting frequency dependence of the out-of-plane susceptibility at the real frequency axis, obtained by analytic continuation. At and close to half filling we find the peak at low frequencies corresponding to a vanishing or small spin gap. However, apart from that, we observe generally two other peaks. One can see that the position of the lower of the two peaks coincides with the low-energy edge of the inter band spectral function $)A_{{\mathbf{Q}},\rm ib}(\omega)=\sum_{\mathbf{k}}\delta(E^{+}_{\mathbf{k}+\mathbf{Q}}-E^{-}_{\mathbf{k}}-\omega)(f(E^{-}_{\mathbf{k}})-f(E^{+}_{\mathbf{k}+\mathbf{Q}}$)), shown in Fig. 1. The low-energy edge of $A_{{\mathbf{Q}},\rm ib}$ is determined by an indirect gap in the electronic spectrum. This peak can be therefore identified with the Higgs mode Sachdev1 ; Sachdev2 ; Sachdev ; SachdevHiggs1 . On approaching the paramagnetic mean-field solution (i.e. at substantial doping levels) this peak shifts to very low energies, and therefore, becomes suppressed by the spin gap. Therefore, at sufficiently large dopings, we observe a single peak in the spin spectral functions. These peaks in the spin spectral function remind us also of the gapped magnetic excitations, discussed recently for the square-lattice Heisenberg model within the bond-operator technique Syromyatnikov .
@@
Figure 3: The correlation length $\xi$ (in logarithmic scale) as a function of hole doping x=1nx=1-n for the temperature T=0.05t{T}=0.05t. Red circles represent data obtained using the full dynamical temporal susceptibility $\chi_{\omega_{n}}$, while green stars correspond to the static approximation applied at all finite bosonic frequencies and blue triangles correspond to the usage of only zero frequency $\omega_{n}=0$. The dashed vertical line indicates electron doping at which the phase transition between spiral and antiferromagnetic magnetic orders occurs. Black diamonds and their black dash-dotted approximation represent the experimental data ExpInel1 ; ExpInel2 on La2-xSrxRuO4 compound, expressed in units of the lattice constant $a=$5.36\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$, and shifted horizontally to adjust the point of AFM to incommensurate state transition.
@@
In summary, we have analyzed the effect of the frequency dependence of temporal spin stiffness on magnetic order in the doped Hubbard model. We find that within the mean-field approach this frequency dependence is crucial to obtain short range magnetic order with a sufficiently small correlation length. In particular, the gapped energy modes, one of which corresponds to the Higgs mode, carry substantial spectral weight, which yields the decrease of correlation length, determined by the sum rule. As a result, we find long-range magnetic order only in the narrow domain of dopings $x\simeq 0.1$, in agreement with the experimental data. We expect that this region of long-range order may be further shrunk by fluctuations, and/or partly unstable with respect to the non-uniform patterns, see, e.g., recent study of Ref. MetznerFiniteLattice, , which finds stripe instability in approximately the same doping range.
@@
Note added. Recently, we became aware of Ref. BonettiNew, , which considered the temperature- and concentration dependencies of spin stiffnesses and correlation length. However, only static temporal spin stiffness was considered in that study.
Acknowledgements. The calculation of correlation length and phase diagram is supported by the Russian Science Foundation (Grant No. 24-12-00186). I. A. G. also acknowledges financial support of the derivation of equations for correlation length from BASIS Foundation (Grant No. 24-1-5-152-1).
@@
1 Connection to the Nonlinear Sigma Model
In order to describe the long-wavelength magnetic fluctuations on the top of the state with long-ranged magnetic order in chargon sector, one can start from the general action of itinerant electrons on the lattice (described, e.g., by the Hubbard model, Eq. (8) of the main text)
$S[c,c^{+}]=\sum_{ij}\int\limits_{0}^{\beta}d\tau\;c_{i}^{+}\left[\left(\frac{\partial}{\partial\tau}-\mu\right)\delta_{ij}-t_{ij}\exp\left(-\vec{r}_{ji}\nabla\right)\right]c_{i}+\int_{0}^{\beta}d\tau H_{\rm int}[c,c^{+}],$ (S1)
where $H_{\rm int}[c,c^{+}]$ is SU(2) symmetric local interaction. Following Refs. BonettiMetzner, ; OurSecond, of the main text, we split the initial fermionic degrees of freedom, described by Grassmann variables $c_{i},c^{+}_{i}$, into the “chargon” degrees of freedom $\psi_{i},\psi^{+}_{i}$, exhibiting long-range magnetic order, and long-wavelength $SU(2)$ rotation field $R$, which introduces spin background distortions at long distances and destroys the long-ranged order, as
$c_{i}$ $={R}_{i}\psi_{i},\,c^{+}_{i}=\psi_{i}^{+}{R}_{i}.$ (S2)
Introducing the gauge field $A_{i\mu}$, defined as
$A_{i\mu}={\rm i}{R}_{i}^{+}\partial_{\mu}R_{i}$ (S3)
with $\partial_{\mu}=(\partial_{\tau},\vec{\nabla})$, the action takes the form
$S[\psi,\psi^{+},A]=\sum_{ij}\int\limits_{0}^{\beta}d\tau\;\psi_{i}^{+}\left[\left(D_{\tau,i}-\mu\right)\delta_{ij}-t_{ij}\exp(-\vec{r}_{ji}\vec{D}_{i})\right]\psi_{i}+\int_{0}^{\beta}d\tau H_{\rm int}[\psi,\psi^{+}],$ (S4)
where $D_{\mu,i}=\partial_{\mu}-{\mathrm{i}}A_{\mu,i}$. Integrating out fermionic degrees of freedom yields the action of the non-linear sigma model (see, e.g., Ref. BonettiMetzner, of the main text)
$S^{\text{eff}}[A]=\frac{1}{2}\iint dx\,dx^{\prime}\,\mathcal{J}^{ab}_{\mu,x;\nu,x^{\prime}}A^{a}_{\mu,x}A^{b}_{\nu,x^{\prime}},$ (S5)
where the components of $A$-field are defined as
$A^{a}_{\mu}=\text{Tr}\left[\sigma^{a}A_{\mu}\right],$ (S6)
and $J_{\mu,x;\nu_{x}}$ is the spin-stiffness tensor, defined by second derivatives of the effective action. To fix the gauge, the infinitesimally small magnetic field should be added (see Refs. BonettiErratum, ; OurSecond, of the main text).
The nonlinearity of the model (S5) appears since the gauge fields $A_{\mu,x}$ are subject to the constraint (S3) with $R_{x}^{+}R_{x}=1$. The constraint can be automatically satisfied if we express the gauge $A$-field in terms of spinon $z,z^{+}$ degrees of freedom (see Ref. BonettiMetzner, of the main text) as
$R=\begin{pmatrix}z_{\uparrow,x}&-z_{\downarrow,x}^{*}\\ z_{\downarrow,x}&z_{\uparrow,x}^{*}\end{pmatrix}R_{0},$ (S7)
with the constraint $z^{+}z=1$; $R_{0}$ represents an arbitrary rotation matrix. For the case of spiral magnetic order we choose $R_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}e^{{\rm i}\pi/4}&e^{{\rm i}\pi/4}\\ -e^{-{\rm i}\pi/4}&e^{-{\rm i}\pi/4}\end{pmatrix}$. In this representation the gauge field $A$ is expressed in terms of $z$-fields as
$A^{x}_{\mu}$ $=2\text{Re}(\tilde{z}^{+}\partial_{\mu}z),$
$A^{y}_{\mu}$ $=2{\rm i}z^{+}\partial_{\mu}z,$
$A^{z}_{\mu}$ $=-2\text{Im}(\tilde{z}^{+}\partial_{\mu}z),$ (S8)
(we have also introduced ”orthogonal” field $\tilde{z}={\rm i}\sigma^{y}z^{*}$).
After substitution of these expressions, the action for interacting $z$-field can be written in the form
$S[z]=2\iint dx\,dx^{\prime}\,\left[J_{\mu x;\nu x^{\prime}}^{\text{op}}(\partial_{\mu,x}z^{+}_{x})(\partial_{\nu,x^{\prime}}z_{x^{\prime}})-J_{\mu x;\nu x^{\prime}}^{\text{op}}(\partial_{\mu,x}z^{+}_{x})P_{x,x^{\prime}}(\partial_{\nu,{x^{\prime}}}z_{x^{\prime}})+J_{\mu x;\nu{x^{\prime}}}^{\text{in}}\,j_{\mu x}j_{\nu{x^{\prime}}}\right]$ (S9)
where $P_{x,x^{\prime}}=\hat{1}-\tilde{z}_{x}\otimes\tilde{z}_{x^{\prime}}^{+}$ ($\hat{1}$ is the unity $2\times 2$ matrix, $\otimes$ corresponds to the direct product), $j_{\mu,x}=-\frac{\mathbf{i}}{2}\left(\partial_{\mu,x}z^{+}_{x}z_{x}-z^{+}_{x}\partial_{\mu,x}z_{x}\right)$, $J^{\text{op}}=J^{xx,zz}$ corresponds to the spin stiffness of the out-plane mode, while $J^{\text{ip}}=J^{yy}$ is the in-plane spin stiffness. The action (S9) can be extended to the action for $N$-component spinon fields. In the limit $N\to\infty$ due to the completeness relation
$z_{x}\otimes z^{+}_{x}+\tilde{z}_{x}\otimes\tilde{z}^{+}_{x}=\hat{1}$ (S10)
one can neglect all the terms except for the quadratic part of the action, which results in the standard N-component spinon action. The constraint $z^{+}z=1$ yields the equation (2) of the main text with the out-of-plane magnetic susceptibility $\chi^{yy}_{\mathbf{q}+\mathbf{Q},\omega_{n}}$ defined in Eq. (1) of the main text, determined by the saddle-point approximation for the resulting bosonic action.
2 q-integration in the sumrule (2) and the choice of the cutoff parameter